### v1

# Prominence-Aware Artifact Detection and Dataset for Image Super-Resolution

Abstract
Generative image super-resolution (SR) is rapidly advancing in visual quality and detail restoration. As the capacity of SR models expands, however, so does their tendency to produce artifacts: incorrect, visually disturbing details that reduce perceived quality. Crucially, their perceptual impact varies: some artifacts are barely noticeable while others strongly degrade the image. We argue that artifacts should be characterized by their prominence to human observers rather than treated as uniform binary defects. Motivated by this, we present a novel dataset of 1302 artifact examples from 11 contemporary image-SR methods, where each artifact is paired with a crowdsourced prominence score. Building on this dataset, we train a lightweight regressor that produces spatial prominence heatmaps and outperforms existing methods at detecting prominent artifacts. We release the dataset and code to facilitate prominence-aware evaluation and mitigation of SR artifacts.
1 Introduction
Single-image super-resolution (SISR) aims to reconstruct high-resolution (HR) images from low-resolution (LR) inputs and has become a cornerstone of low-level vision tasks. Recent progress with deep learning and generative adversarial networks (GANs) has greatly improved perceptual quality but also introduced a critical challenge: generation of visually unpleasant artifacts. These artifacts—usually unnatural patterns, smeared faces, and texture distortions—degrade the perceived quality of SISR outputs and hinder adoption. Even the latest advancements, such as transformer- and diffusion-based methods (Liang et al., 2021; Yu et al., 2024), remain prone to creating artifacts.
Despite SISR’s growing popularity, research on detecting SR artifacts remains scarce. LDL (Liang et al., 2022) and DeSRA (Xie et al., 2023a) both rely on residual statistics to localize artifacts but differ in supervision: LDL uses HR references and regularizes SR models during training, whereas DeSRA contrasts outputs from the same backbone trained with GAN vs. MSE losses and enables few-shot fine-tuning of existing SR models. Approaches such as PAL4VST (Zhang et al., 2023) cast artifact detection as a segmentation problem by predicting a binary mask from the output image.
These methods rely on manually annotated datasets that contain binary artifact masks. We argue that this limitation is critical: artifacts vary in their prominence to viewers. For example, distortions to regular structures such as buildings, or to recognizable objects such as human faces, easily draw attention and can be distressing to viewers (Figure 1). On the other hand, artifacts on water, grass, and other organic matter can be almost unnoticeable (Figure 2). Treating these different cases as equal carries the risk of overfitting a detection method to less important artifacts while missing the disturbing ones, thus degrading the viewing experience.
To address this limitation, we created a comprehensive dataset of 1302 SR artifact examples generated by 11 contemporary SISR methods from 500 source images, each annotated with a prominence score from extensive crowdsource assessments. We further collected prominence scores for all 593 artifact examples in the DeSRA dataset (originally annotated in lab by Xie et al.) and found that nearly half of these artifacts aren’t prominent to most viewers.
Building on our dataset, we propose a prominence-modeling method to detect and quantify SISR artifacts. Our study evaluated existing artifact-detection and image-quality metrics for their ability to predict prominence and trained a lightweight regressor to map their outputs to spatial prominence heatmaps. In parallel, we adapted full-reference methods using the output of RLFN (Kong et al., 2022)—a real-time SR method robust to artifacts—as pseudo ground truth (pseudo-GT), showing that this does not significantly degrade detection quality. This makes our approach, and full-reference baselines, applicable in real-world scenarios where high-resolution ground truth is unavailable.
Figure 1: Examples of prominent artifacts detected by the proposed method. Top: example from our proposed dataset. GFPGAN failed to restore holes on the radio panel. Bottom: example from the DeSRA dataset. LDL produced an irregular line pattern on the carpet. Prominence denotes the percentage of annotators confirming the artifact in the highlighted area (Section 3).
Figure 2: Non-prominent artifacts detected by existing methods. Top: example from our proposed dataset. The water surface was incorrectly restored by GFPGAN producing an artifact, but it’s a natural surface, so this artifact is not prominent to humans. Bottom: example from the DeSRA dataset. SwinIR produced a texture artifact (vertical lines) on the floor, but it’s in a non-salient region, so it’s not prominent to humans. Our method correctly marks both examples with low prominence.
Our main contributions are the following:
  1. 1.
    We present the first dataset of its kind, containing 1302 SISR artifact examples with crowdsourced prominence annotations; it features diverse distortions from 11 SISR methods. To adapt for visual inspection the tight masks that artifact-detection methods produce, we propose a simple mask-postprocessing algorithm. We additionally collected prominence annotations for the 593 artifact examples in the DeSRA dataset (Xie et al., 2023a).
  2. 2.
    We propose a perceptual-prominence-modeling method that detects and quantifies artifacts in SR images and that outperforms existing approaches, judging by thorough objective and subjective evaluation. We show that our method, relative to prior art, is more applicable to fine-tuning of SR models for artifact reduction. Furthermore, we demonstrate that real-time SR can be a useful pseudo-GT for full-reference metrics.
  3. 3.
    Using our proposed artifact-detection method and existing alternatives, along with our crowdsourced prominence annotation methodology, we analyzed 11 SR methods for their propensity to generate artifacts. We show that even the newest high-quality methods such as SUPIR (Yu et al., 2024) are highly susceptible to this problem.
We published our code and datasets with prominence annotations at tinyurl.com/2u9zxtyh.
2 Related work
Single-image super-resolution (SISR) has undergone extensive study over the past decade. Early methods optimized pixel-wise losses such as $L_{1}$ and $L_{2}$, which improved PSNR/SSIM but failed to recover realistic details. GAN-based approaches improved perceptual sharpness and introduced realistic degradation pipelines for training (Ledig et al., 2017; Zhang et al., 2021; Wang et al., 2021b).
Recent SISR developments span Transformer-based architectures, diffusion-driven models, and hybrid generative priors. Transformer approaches such as HAT (Chen et al., 2023) and DRCT (Hsu et al., 2024b) use multiscale attention and dense-residual connections to enhance quality while remaining efficient. Diffusion-based methods like ResShift (Yue et al., 2023) and SinSR (Wang et al., 2024b) accelerate sampling via residual shifting and deterministic distillation, while StableSR (Wang et al., 2024a) leverages Stable Diffusion (Rombach et al., 2021) priors for real-world upscaling and SUPIR (Yu et al., 2024) scales models and data for text-guided photo-realistic restoration. However, these advances bring challenges, particularly the introduction of visual artifacts.
SISR evaluation has traditionally relied on full-reference metrics such as PSNR and SSIM (Wang et al., 2004), which assess reconstruction fidelity but correlate poorly with perceptual quality—especially for GAN-based outputs where details and artifacts are entangled. No-reference and perceptual metrics such as NRQM (Ma et al., 2017), LPIPS (Zhang et al., 2018), and DISTS (Ding et al., 2022) were introduced to better align with human perception and are now widely adopted in SR benchmarks. Some techniques also aim to make metrics more resistant to artifacts; for example, ERQA (Kirillova. et al., 2022) evaluates detail restoration by matching edges in reference and test images. However, in practice existing quality metrics are not sensitive enough to artifacts.
Detection and mitigation of SISR artifacts has garnered increasing attention because these artifacts reduce perceptual quality. LDL (Liang et al., 2022) predicts pixel-level artifact maps from local residual statistics. Xie et al. (2023a; b) introduced an in-lab annotated dataset with binary SR artifact masks and, building on it, proposed DeSRA that contrasts GAN-SR and MSE-SR outputs to identify artifact-prone regions, then fine-tunes the SR model on a few samples to suppress those regions.
A complementary line of work treats artifact detection as segmentation, training networks on datasets with pixel-level defect maps. Given only an input image, these models predict an artifact mask. Approaches such as PAL4Inpainting (Zhang et al., 2022) and PAL4VST (Zhang et al., 2023) show strong generalization across generative vision tasks by localizing perceptual artifacts.
Concurrently, Ren et al. (2025) propose Hallucination Score (HS), which queries a multimodal LLM to assign a single image-level hallucination rating for SR outputs, showing strong alignment with human judgments. The main drawback of this approach is that it lacks spatial localization, which is critical for downstream tasks such as artifact mitigation, SR model fine-tuning, and for handling cases where different regions of an image exhibit different types of artifacts.
Despite these advances, the shortage of annotated datasets that explicitly focus on SR artifacts limits most approaches. While datasets such as DeSRA provide binary artifact masks, most methods remain constrained by the absence of richer annotations (e.g., prominence levels), which limits their generalization and robustness in real-world scenarios. Our work addresses this shortfall by introducing a novel dataset annotated with artifact regions and prominence scores, along with a prominence-aware detection method that supports SR fine-tuning and reveals that even the latest models like SUPIR remain highly artifact-prone.
3 Artifact dataset
Existing datasets such as DeSRA (Xie et al., 2023a) contain only binary artifact masks, without information on how noticeable the artifacts are to viewers. To enable research on artifact prominence, we introduce a dataset of 1302 artifact examples, each annotated with both a binary mask and a prominence score derived from crowdsourced assessments.
Our dataset is based on Open Images (Kuznetsova et al., 2020), a diverse collection of about 300,000 natural images (CC BY 2.0). We randomly selected 2,101 source photos, each 768×1,024 pixels. These photos then underwent 4× bicubic downsampling followed by upsampling with 11 popular SR methods (Yu et al., 2024; Wang et al., 2021b; Chen et al., 2023; Hsu et al., 2024b; Yue et al., 2023; Wang et al., 2024b; a; Cai et al., 2019; Liang et al., 2021; Wang et al., 2021a; Wu et al., 2024), yielding 23,111 images for artifact search. We obtained the initial binary artifact masks through manual annotation, and by running existing visual-quality metrics (SSIM, DISTS, LPIPS) and artifact-detection algorithms (LDL, DeSRA, in-progress versions of our method). For each of these algorithms we selected the 100 strongest detected distortions and manually discarded images without artifacts. The remaining images underwent crowdsourced prominence annotation, resulting in 697 artifact examples. The evaluation in Section 5.3 yielded 605 more examples.
We additionally collected prominence annotations for all 593 images from the DeSRA artifact dataset.
Figure 3: Example of our postprocessing technique increasing artifact visibility. On the left is the original mask from the DeSRA dataset; on the right is the mask after postprocessing.
Figure 4: Prominence distribution for our dataset (based on Open Images) and for the DeSRA dataset (for which we collected prominence annotations).
Figure 5: Interface presented to the viewers during subjective data collection.
Table 1: Comparison between the DeSRA and our proposed artifact dataset.
DeSRA Ours
# SR methods 3 11
# Artifacts 593 1302
Mask source in-lab automatic
Label type binary prominence
3.1 Crowdsourced annotation setup
We used Toloka.ai to crowdsource the data collection. Participants view pairs of images labeled “Original” and “Upscaled,” with the artifact region visually highlighted. We ask them whether the highlighted region contains a distorted object or texture. Figure 5 shows an example question.
Every image receives a ranking by 30 different participants. We compute prominence as the proportion of votes indicating the artifact is present. Before receiving access to the main questions, participants must answer four training questions, for which the correct answers are explained, followed by four test questions with hidden correct answer. Afterward, to ensure integrity, every group of 20 questions contains 4 random control questions. All responses from participants who mistakenly answer any control question are discarded. In total, 264 participants successfully completed the annotation. Appendix A analyses the impact of the participant count on the answer variability.
3.2 Mask postprocessing
An artifact-detection method aims to output a tight mask around an artifact, since doing so is more useful for further analysis and for downstream tasks such as automatic correction. But tight masks make it harder to visually judge whether the masked area contains an artifact.
To remedy this, we apply morphological operations to the masks before showing them to participants:
  1. 1.
    Open with a 25×25 square kernel to remove small dots in the mask.
  2. 2.
    Dilate with a 64×64 circular kernel so the mask includes context around an artifact.
  3. 3.
    Close with a 25×25 square kernel to eliminate holes and step away from the image borders.
The example in Figure 4 shows how a tight mask makes an artifact harder to notice compared with our postprocessing result. We verified the validity of this postprocessing step by running the crowdsourced prominence annotation twice on the DeSRA dataset: once with our postprocessing and once with unmodified masks. For this comparison, we ensured that separate groups of participants conducted the annotations, all while preserving the exact question order. As a result, the average artifact prominence for the entire DeSRA dataset was 49.4% with our postprocessing and 47.7% with the original masks, indicating the postprocessing helps viewers judge an artifact’s presence.
3.3 Dataset analysis
Table 1 and Figure 4 compare our dataset to DeSRA and show prominence distributions.
Our dataset has a bias toward zero prominence (no or unnoticeable artifacts); the reason is that we seeded the masks with results from existing image-quality metrics poorly adapted to finding artifacts and from existing artifact-detection methods that lack the ability to differentiate between barely visible and highly visible artifacts. This bias has implications for objective evaluation, potentially skewing results, and can induce overfitting during model training—a concern we address in Section H.4.
The DeSRA dataset has a more even split, with a tendency toward middle-ground artifact prominence. Notably, even though artifacts in DeSRA underwent annotations in lab, almost half of them have a prominence below 50%—that is, most viewers fail to notice them. This result confirms that binary masks are insufficient for accurate SR-artifact evaluation.
Figure 6: Architecture of the proposed artifact prominence metric. The input image is upscaled by the target SR and by RLFN as described in Section 4.3. Then, we compute three features described in Section 4.1. Finally, we run our prominence regressor described in Section 4.2.
4 Artifact prominence metric
Our goal is to predict a spatial heatmap of artifact prominence for a super-resolved image, given the corresponding low-resolution input. In this heatmap, higher values indicate regions where artifacts are more severe (i.e., more prominent to human observers).
Figure 6 shows an outline of our proposed method. Our method aggregates three features chosen from existing visual-quality metrics and artifact-detection methods. On the basis of these features, we trained a multilayer-perceptron regressor. The resulting model finds prominent image artifacts more efficiently compared with any of these features individually and compared with other approaches.
4.1 Input features
We selected features based on their proven performance for evaluating and detecting texture distortions. So, these features estimate not only visual quality of images, but also take into account structural similarity of a high-resolution image and an upscaled image.
The first feature is the visual-image-quality metric DISTS (Ding et al., 2022), which takes into account possible texture distortions and their impact on visual perception. Since DISTS is trained on natural images, it is good at detecting unnatural degradations such as SR-produced artifacts. DISTS produces a single score for a given image, so we applied it block-wise to obtain a feature map. We computed DISTS in 16×16-pixel blocks, the metric’s minimum input size.
The second feature, which we call ssm_jup, is adapted from the small-color-artifact detector from Tsereh et al. (2024), itself based on LDL (Liang et al., 2022). It targets small-scale image distortions and was shown to be effective for finding JPEG AI compression artifacts. Unlike Tsereh et al. (2024), we use all RGB channels rather than only chromatic U and V components to capture texture distortions. As with LDL, this feature requires a reference image upscaled by a more artifact-resistant method; we chose bicubic interpolation for this reference input.
The last feature, bd_jup, is the weighted sum of LPIPS (Zhang et al., 2018) and ERQA (Kirillova. et al., 2022) applied block-wise. LPIPS measures how well the upscaled image preserves perceptual quality, and is widely used in SR evaluation. Meanwhile, ERQA is designed to assess the preservation of object details and boundaries. For LPIPS, we used 32×32-pixel blocks with stride of 16. ERQA uses 8×8 blocks with no overlap. LPIPS is weighted 3:2 compared with ERQA.
4.2 Prominence regressor
We aggregated the features described above using a shallow multilayer perceptron (MLP). This network takes as input the feature values, passes them through three fully connected layers with ReLU activations, and outputs a single prominence value. The MLP independently processes each pixel of the input-feature heatmaps, predicting an artifact’s prominence in that pixel. Since our selected features employ the pixel’s neighborhood as well as wider input-image context, our final method also considers the wider context despite the MLP’s pixel-wise application.
We also experimented with CNN- and tree-based models, but the shallow MLP achieved the best overall performance (Sections H.6 and H.7).
4.3 Adapting full-resolution metrics with real-time-SR pseudo-GT
Full-reference metrics provide more-accurate detail-restoration-quality scores for SR by employing pixel-level information from the reference image. The use of such metrics in SR creates difficulties, however, since the SR-output resolution is higher than that of the original low-resolution frame.
To employ full-reference metrics, we propose the following pseudo-GT pipeline. We applied a lightweight real-time SR method to the original low-resolution frame, thereby obtaining a pseudo-GT, and then calculated the metric for this pseudo-GT and the SR output. We noticed that real-time SR methods, such as SPAN (Hu et al., 2025) and RLFN (Kong et al., 2022), produce outputs that, despite trailing heavier SR models in visual quality, are devoid of major visual artifacts. When serving as pseudo-GT for full-reference metrics, the resulting artifact-detection performance drop is small compared with using the original HR frames, as Section 5.2 shows. This approach enables our method to serve in real-world upscaling where the high-resolution GT frames are unavailable.
4.4 Training
We train our MLP prominence regressor using Adam on a training subset of 374 artifact examples from our dataset described in Section 3. The MLP predicts a prominence value for each pixel of the input image. We compute the mean predicted prominence inside and outside the binary artifact mask from the dataset. The training loss consists of two $L_{2}$ components:
$\mathcal{L}=L_{2}(\mathrm{MeanInside},\mathrm{GT\ Prominence})+L_{2}(\mathrm{MeanOutside},0).$ (1)
The model is trained to predict the ground-truth prominence value inside the binary mask, and 0 (no artifact) outside it. Thanks to small model size, the training converges quickly, usually in around 10–30 epochs. One training epoch takes about 13 seconds on an Nvidia RTX 3090 GPU.
Figure 7: Overview of our objective evaluation pipeline. Metric heatmaps are binarized and compared with GT artifact masks to obtain TP/FP/FN/TN regions. Raw heatmap values within these regions are then used to compute soft precision and recall, taking into account crowdsourced artifact prominence.
5 Experiments
Our evaluation comprises three components. Section 5.1 provides an overview of our approach to objectively evaluating artifact-detection methods using our prominence dataset, then Section 5.2 compares our method to existing work with objective scores. Next, Section 5.3 evaluates methods on the primary downstream task: finding artifacts prominent to human viewers. Finally, Section 5.4 evaluates methods on the secondary downstream task: reducing SR proneness to artifacts.
5.1 Objective evaluation methodology
Our evaluation follows standard binary classification methodology—binarizing detection outputs with thresholds (selected for each method in evaluation) and computing TP, FP, FN—but modifies the precision and recall to weight detections by their prominence scores (Figure 7). Following Rachakonda & Bhatnagar (2021); Harju & Mesaros (2023), we treat artifact labels as graded values rather than binary. In our setting, each label is a prominence score rather than a probability of class membership. A low prominence score indicates that the masked area contains no artifact, so a positive detection in such cases should be penalized. We implement this using a margin $\kappa=0.3$, yielding:
$Prec^{pr}=\frac{\sum_{i\in\text{images}}TP_{i}\ast(\text{p}_{i}-\kappa)}{\sum_{i\in\text{images}}(TP_{i}+FP_{i})};\quad Rec^{pr}=\frac{\sum_{i\in\text{images}}TP_{i}\ast(\text{p}_{i}-\kappa)}{\sum_{i\in\text{images}}(TP_{i}+FN_{i})}.$ (2)
We set $\kappa$ = 0.3 to reward metrics that detect even less prominent artifacts: if more than 30% of viewers reported seeing the artifact, we already already consider it prominent enough to detect. This choice was guided by dataset statistics and observed viewer consistency.
From $Prec^{pr}$ and $Rec^{pr}$, we compute F1-score and PR-AUC. Note that unlike binary classification, scores computed this way have a different range from $[0,1]$ (see Appendix F). However, they remain suitable for ranking methods relatively to each other.
Table 2: Results on the full proposed and DeSRA datasets. PR-AUC aggregates over thresholds so it is shown once. The last column shows the method’s average ranking by PR-AUC compared to others.
Proposed Dataset DeSRA Dataset Avg.
Reference Input Original HR SPAN RLFN MSE-SR SPAN RLFN Rank
Method F1-score PR-AUC F1-score PR-AUC F1-score PR-AUC F1-score PR-AUC F1-score PR-AUC F1-score PR-AUC PR-AUC
LDL (t=0.005) 0.0275 0.0039 0.0197 0.0034 0.0035 0.0008 0.1670 0.0518 0.1618 0.0486 0.1622 0.0503 4.7
[…]
ERQA (t=0.55) 0.0028 0.0001 -0.0137 -0.0002 -0.0195 -0.0004 0.0396 0.0026 0.0474 0.0017 0.0399 0.0025 9.0
PAL4Inpaint (bin., no-ref) 0.0117 N/A 0.0117 N/A 0.0117 N/A 0.0609 N/A 0.0609 N/A 0.0609 N/A N/A
PAL4VST (bin., no-ref) 0.0062 N/A 0.0062 N/A 0.0062 N/A 0.0054 N/A 0.0054 N/A 0.0054 N/A N/A
[…]
bd_jup (t=0.1) 0.0043 0.0027 0.0105 0.0013 0.0074 0.0017 0.1175 0.0230 0.0920 0.0135 0.1181 0.0244 6.9
ssm_jup (t=0.2) 0.0251 0.0012 0.0144 0.0011 0.0180 0.0009 0.1769 0.0377 0.1426 0.0251 0.1690 0.0346 6.4
[…]
Table 3: Results on the prominent subset of the proposed and DeSRA datasets.
Proposed Dataset DeSRA Dataset Avg.
Reference Input Original HR SPAN RLFN MSE-SR SPAN RLFN Rank
Method IoU PR-AUC IoU PR-AUC IoU PR-AUC IoU PR-AUC IoU PR-AUC IoU PR-AUC PR-AUC
LDL (t=0.005) 0.1043 0.1387 0.1788 0.3110 0.1779 0.3361 0.3724 0.4687 0.3896 0.3418 0.3443 0.4182 6.0
[…]
ERQA (t=0.55) 0.2495 0.1220 0.1670 0.1052 0.1492 0.1052 0.0523 0.0100 0.0684 0.0154 0.0514 0.0087 10.0
PAL4Inpaint (bin., no-ref) 0.0753 N/A 0.0753 N/A 0.0753 N/A 0.1139 N/A 0.1139 N/A 0.1139 N/A N/A
PAL4VST (bin., no-ref) 0.0463 N/A 0.0463 N/A 0.0463 N/A 0.0140 N/A 0.0140 N/A 0.0140 N/A N/A
[…]
bd_jup (t=0.1) 0.2843 0.3311 0.2475 0.2342 0.2434 0.2275 0.3580 0.1609 0.2798 0.1221 0.3625 0.1773 7.9
ssm_jup (t=0.2) 0.2368 0.2127 0.2133 0.2273 0.2221 0.2411 0.4032 0.3889 0.3770 0.2737 0.3930 0.3646 7.5
DeSRA (t=0.3) 0.2560 0.3173 0.1296 0.3358 0.1205 0.3025 0.5277 0.6928 0.3707 0.2910 0.5082 0.6614 3.8
[…]
5.2 Objective prominence-metric evaluation
We evaluate our method using both our proposed methodology, and the methodology from DeSRA (Xie et al., 2023a), to avoid bias towards our own dataset and scoring. Following DeSRA, we selected binarization thresholds for all methods by maximizing the Precision × Recall product on the prominent subset of the DeSRA dataset. The thresholds are shown in all tables as t=0.xx.
Table 2 compares our method with other approaches on the basis of our prominence datasets, under different choices of reference input as described in Section 4.3. These include original high-resolution frames, the lightweight SR models SPAN (Hu et al., 2025) and RLFN (Kong et al., 2022), and, for DeSRA where HR frames are unavailable, the authors’ MSE-SR model trained with an MSE loss. Since the MSE-SR weights were not released, this reference cannot be used in other experiments.
Our method delivers the most consistent performance across all experiment settings, as reflected by its best average ranking in PR-AUC. These results also confirm that using pseudo-GT remains practical for artifact detection. We adopt RLFN pseudo-GT for all other experiments in the paper.
Next, we followed the DeSRA comparison methodology and computed binary precision, recall, and intersection-over-union (IoU) on the same datasets (no $\kappa$ margin). This comparison, however, only considered a subset with prominence values above 50%; doing so yields a more accurate evaluation, as artifacts with low prominence values are barely noticeable and should be considered false positives. Table 3 shows the results. Our method outperforms competitors in most experiment settings.
We also evaluated our method on a learning-based compression task, as described in Appendix G. You can find $Prec^{pr}$ and $Rec^{pr}$ values for all methods in Section H.8.
5.3 Evaluating robustness to SR artifacts
We evaluated the robustness of 11 popular SR models to generating artifacts, following the preparation process from Section 3, but with no manual mask curation. For each metric, we selected the 10 strongest artifacts per SR, yielding 653 in total. We then collected prominence values for these masks via crowdsourcing. Our results report the total number of artifact masks that the metrics produced, their mean prominence, and the number of “confident” masks corresponding to highly visible artifacts (50% prominence or higher).
Table 4: Crowd-sourced prominence results across SR models.
SR Type Masks Found Mean Prominence Conf. Masks Found
DRCT (Hsu et al., 2024a) Transformer 43 11.85% 1
HAT-L (Chen et al., 2023) Transformer 53 13.53% 1
SinSR (Wang et al., 2024b) Diffusion 60 19.39% 3
ResShift (Yue et al., 2023) Diffusion 60 20.22% 7
StableSR (Wang et al., 2024a) Diffusion 60 28.55% 14
RealSR (Cai et al., 2019) CNN 51 28.70% 10
SeeSR (Wu et al., 2024) Diffusion 61 32.34% 14
GFPGAN (Wang et al., 2021a) CNN 45 32.74% 11
SwinIR (Liang et al., 2021) Transformer 59 41.09% 17
SUPIR (Yu et al., 2024) Diffusion 70 45.29% 20
RealESRGAN (Wang et al., 2021b) CNN 61 48.42% 19
Table 5: Crowd-sourced prominence results across artifact detection methods.
Method Masks Found Mean Prominence Conf. Masks Found Comb. Score
LDL (t=0.005) 12 77.11% 11 8.48%
bd_jup (t=0.1) 110 17.03% 13 2.21%
LDL (t=0.0005) 51 34.84% 15 5.23%
LDL (t=0.001) 40 43.00% 16 6.88%
ssm_jup (t=0.15) 110 23.09% 20 4.62%
SSIM (t=0.55) 74 36.62% 26 9.52%
DeSRA 110 32.03% 31 9.93%
DISTS (t=0.25) 108 38.80% 38 14.75%
Ours (t=0.3) 99 41.25% 38 15.67%
Table 5 shows the results grouped by SR model. DRCT (Hsu et al., 2024a) and HAT-L (Chen et al., 2023), both Transformer based, show excellent results with nearly zero prominent artifacts detected. Next are three diffusion-based methods: SinSR (Wang et al., 2024b), ResShift (Yue et al., 2023), and StableSR (Wang et al., 2024a). Then, Table 5 shows the results grouped by artifact-detection method. Ours shares first place with DISTS in number of confident masks found while beating it in mean artifact prominence. DeSRA ranked third in this evaluation.
Also note that LDL with a 0.005 threshold finds regions with highly visible artifacts, but it trails far behind other methods in total number of confident masks. To account for both of these scores, we multiplied them analogously to the precision × recall product of Xie et al. (2023a); the results are in the last column. We tested LDL at two lower thresholds, which increased the total masks found, but they mainly captured non-prominent artifacts, yielding worse combined score.
Appendix E contains extra crowd-sourced evaluation results on other SR datasets.
Table 6: Results of fine-tuning SR models on artificial GT constructed with different methods.
Target SR LDL Liang et al. (2022) RealESRGAN Wang et al. (2021b) SwinIR Liang et al. (2021) Avg.
[…]
DISTS (t=0.25) 27.00 33.51 11.70 33.47 32.66 15.58 15.23 26.47 13.73 3.3
LDL (t=0.005) 17.26 90.43 6.91 18.32 88.94 11.56 4.22 77.94 21.08 4.9
LPIPS (t=0.25) 25.18 15.43 19.68 33.66 8.54 34.17 11.21 2.94 53.43 2.2
ERQA (t=0.55) 0.67 100.00 0.00 1.22 99.50 0.00 0.19 98.04 0.00 6.0
DeSRA 29.18 54.26 25.00 33.66 34.17 56.78 8.98 32.84 28.43 2.9
Ours (t=0.3) 34.71 20.74 45.21 38.01 14.57 55.78 11.86 6.37 57.35 1.6
5.4 Fine-tuning SR models to reduce artifacts
We follow the methodology from Xie et al. (2023a) and fine-tune SR models to reduce artifacts using artifact-detection methods. We fine-tune each SR with each method and compare the results.
Fine-tuning optimizes the pixel-wise MSE loss between the SR model’s output and an artificial GT image, constructed by replacing the artifact regions on the SR model’s output with output from RLFN. The artifact regions are the binarized output masks of a given artifact-detection method.
We used the DeSRA dataset for both training and testing, with the split determined by the target SR model (RealESRGAN, LDL, or SwinIR). For a given model, images with an artifact mask for that model were reserved for testing, while the remaining images were used for fine-tuning. This setup enabled us to evaluate IoU, artifact removal, and addition against GT masks on the held-out test set. Each model was trained on roughly 300 images and tested on about 200.
[…]
6 Conclusion
In this work, we address the challenge of visual artifacts in single-image super-resolution, a problem that plagues even the newest, most capable models such as Yu et al. (2024). We argued that artifacts should be characterized by their prominence to human observers rather than by binary masks, as the perceptual impact of artifacts varies significantly. We validated this perspective by showing that most viewers fail to notice almost half of the artifacts in the DeSRA dataset, which was annotated in-lab.
Our primary contribution is a new dataset of 1302 artifact examples with prominence annotations from 11 contemporary SISR methods. We also provide prominence annotations for all 593 artifacts from the DeSRA dataset.
Building on this, we developed an MLP-based method for SR-artifact detection that outputs prominence heatmaps, moving beyond simple binary masks. We conducted broad evaluations showing that compared with other methods, ours can more accurately evaluate the artifact severity of SR models. We showed the practical utility of our method in detecting prominent artifacts and fine-tuning SR models for artifact suppression.
The implications of our work extend beyond SISR. The concept of artifact prominence is likely applicable to other image processing and restoration tasks, as suggested by our preliminary findings on a dataset of JPEG AI artifacts. Furthermore, prominence-aware metrics could guide future SR research to focus on structured regions where artifacts are most visible.
We acknowledge the limitations of our study. The artifact masks in our dataset are approximate, as delineating exact artifact boundaries is ambiguous even for human annotators. Instead, we seed masks using existing methods, introducing some inaccuracy. Additionally, our proposed pseudo-GT approach relying on a lightweight SR model can lead to false positive detections when it fails to reconstruct fine textures that a more powerful model resolves correctly. Finally, the lack of existing prominence datasets limits our ability to faithfully validate our method.
There are several promising directions for future research. A natural extension is towards video super-resolution: while our work applies on a per-frame basis, we do not address temporal artifacts such as flickering. To mitigate the failure mode of a single lightweight SR, using an ensemble of models for pseudo-GT could be explored. Finally, higher-capacity models such as Yu et al. (2024) start to shift from simple texture distortions to more semantic artifacts like object replacement, inviting further investigation.
Our code and datasets, with prominence annotations, are available at tinyurl.com/2u9zxtyh.
Acknowledgments
This work was supported by the The Ministry of Economic Development of the Russian Federation in accordance with the subsidy agreement (agreement identifier  [^00] C313925P4H0002; grant No 139-15-2025-012).
The research was carried out using the MSU-270 supercomputer of Lomonosov Moscow State University.
Reproducibility statement
We have made our code, our artifact prominence dataset, and the prominence annotations for the DeSRA dataset publicly available, linked in Sections 1 and 6. The paper provides a detailed description of the dataset collection and crowdsourced assessment methodology in Section 3, the architecture and features of the proposed artifact detection method in Section 4, and the evaluation methodology in Section 5. Appendix B further describes the metrics used in the SR fine-tuning experiment.
Appendix A Crowdsourced annotation dispersion analysis
Figure 8: Bootstrap-analysis results for an image with a highly prominent artifact (left) and barely prominent artifact (right). Red line indicates our chosen assessor count of 30.
Our crowdsourced prominence-annotation work, described in Section 3.1, involved 30 participants ranking every image separately. This appendix provides our motivation for choosing this number by analyzing the answer dispersion.
We took 11 SR-upscaled images with artifacts of varying intensity and conducted crowdsourced prominence annotation following the same procedure, but with a higher participant count: every image underwent ranking by 264 people. Next, we performed a bootstrap analysis on the votes. For each assessor count $k$ from 1 to 100, the analysis randomly sampled $k$ votes with replacement and computed the prominence from these votes. This procedure repeated nn=1000 times; we then computed 95% confidence intervals for each assessor count $k$.
Figure 8 shows these confidence intervals for two sample images: one with a highly prominent artifact and another with a barely prominent artifact. In cases with few assessors (1–5), the confidence interval frequently spans the whole prominence range from 0% to 100%, meaning any given 5 assessors may all state that an artifact is present or absent. This is especially true for unclear cases at around 50% prominence. As the assessor count grows, the confidence interval shrinks, reaching approximately ±10% at 100 assessors.
For the rest of our annotation process we chose an assessor count of 30 as a reasonable compromise between the confidence of the result (±20%) and the time/cost of using many assessors.
Appendix B SR fine-tuning scoring and dilation details
Figure 9: Examples of SR artifacts removed after fine-tuning with the proposed metric, along with corresponding detected-artifact masks. The DeSRA dataset images are under the ImageNet license.
Figure 10: Schematic view of our SR fine-tuning pipeline.
Table 7: Results of fine-tuning SR models on artificial GT constructed with different methods with and without dilation. The top half of the table matches the table in the main paper text.
Target SR LDL (Liang et al., 2022) RealESRGAN (Wang et al., 2021b) SwinIR (Liang et al., 2021)
[…]
Dilation = True
DISTS (t=0.25) 27.00 33.51 11.70 0.03 93.27 33.47 32.66 15.58 0.03 96.49 15.23 26.47 13.73 0.02 93.38
LDL (t=0.005) 17.26 90.43 6.91 0.25 97.95 18.32 88.94 11.56 0.13 98.40 4.22 77.94 21.08 0.04 98.03
LPIPS (t=0.25) 25.18 15.43 19.68 0.15 82.50 33.66 8.54 34.17 0.01 90.56 11.21 2.94 53.43 0.00 83.55
ERQA (t=0.55) 0.67 100.00 0.00 0.53 35.34 1.22 99.50 0.00 0.56 36.46 0.19 98.04 0.00 0.20 40.95
DeSRA 29.18 54.26 25.00 0.58 71.02 33.66 34.17 56.78 0.25 80.06 8.98 32.84 28.43 0.12 51.69
Ours (t=0.3) 34.71 20.74 45.21 0.03 97.20 38.01 14.57 55.78 0.03 98.09 11.86 6.37 57.35 0.00 91.95
Dilation = False
DISTS (t=0.25) 24.97 52.13 2.13 0.05 88.22 32.93 47.24 7.04 0.04 94.17 14.58 25.98 7.35 0.01 91.03
LDL (t=0.005) 12.47 100.00 0.00 0.36 77.24 15.06 99.50 0.50 0.38 82.11 3.65 98.04 0.98 0.07 79.31
LPIPS (t=0.25) 23.65 18.09 17.02 0.16 78.20 25.13 15.08 32.20 0.04 87.38 10.90 3.92 47.06 0.00 82.39
ERQA (t=0.55) 0.34 98.94 0.00 0.37 19.04 0.45 99.50 0.00 0.51 20.40 0.06 97.06 0.00 0.12 24.18
DeSRA 27.59 57.45 21.81 0.53 68.33 33.23 39.20 50.25 0.31 77.36 8.47 35.29 25.98 0.15 49.89
Ours (t=0.3) 31.63 45.21 12.77 0.14 88.83 36.83 48.24 23.12 0.12 93.49 11.01 24.02 30.88 0.00 87.50
Table 8: Results of fine-tuning SR models on artificial GT constructed with different methods (absolute values).
Target SR LDL (Liang et al., 2022) RealESRGAN (Wang et al., 2021b) SwinIR (Liang et al., 2021)
Method IoU before IoU after PixFrac before PixFrac after IoU before IoU after PixFrac before PixFrac after IoU before IoU after PixFrac before PixFrac after
DISTS (t=0.25) 29.51 2.51 24.69 2.00 34.30 0.83 9.44 0.30 16.53 1.30 20.73 1.92
LDL (t=0.005) 18.08 0.82 8.20 0.47 18.88 0.56 3.10 0.18 4.27 0.05 3.90 0.21
LPIPS (t=0.25) 30.62 5.44 8.48 2.29 36.37 2.71 1.50 0.18 12.20 0.99 5.02 0.96
ERQA (t=0.55) 4.64 3.97 4.95 3.96 5.94 4.72 5.17 3.04 1.00 0.81 2.52 1.73
DeSRA 35.73 6.55 9.04 4.68 35.35 1.69 4.68 0.36 13.06 4.08 5.26 2.73
Ours (t=0.3) 35.86 1.15 12.57 0.99 38.31 0.30 2.28 0.18 12.06 0.20 5.77 0.42
Figure 11: Evolution of metric values over 10 epochs of RealESRGAN fine-tuning with DeSRA compared to the proposed method.
Section 5.4 overviews our fine-tuning process to reduce artifacts of existing SR models. Figure 9 shows fine-tuning examples. Here, we provide more details on the pipeline and on the scoring process.
Figure 10 shows our fine-tuning pipeline. The input low-resolution image is upscaled with the target SR and RLFN models, and the results are passed to the artifact metric. Then, the regions on the target-SR output where the artifacts were detected are replaced with regions from the RLFN output. The resulting artificial GT image is then used as a target to fine-tune the target-SR model.
Table 6 in the main section uses several scores to compare fine-tuning results. Below are the detailed descriptions of these scores. In the following formulas, $A$ represents the set of pixels detected by a metric after fine-tuning, $B$ represents the set of pixels detected by a metric before fine-tuning, and $GT$ represents the set of pixels belonging to the ground-truth artifact mask annotations (from the DeSRA dataset).
  • $\Delta\text{IoU}$. Average reduction of IoU with GT artifact mask: $\frac{\left|B\cap GT\right|}{\left|B\cup GT\right|}-\frac{\left|A\cap GT\right|}{\left|A\cup GT\right|}$.
  • $\text{Rem}_{img}$ and $\text{Add}_{img}$, image-wise removal and addition rates. The ratio of images where $\left(\left|A\cap B\right|=0\right)\wedge\left(\left|B\right|\neq 0\right)$ determines whether the artifact was removed, and the ratio where $A\cup B>B$ determines whether a new artifact was introduced.
  • $\text{Rem}_{img}$: the image-wise removal rate. It represents the percentage of images in test set in which fine-tuning removed artifact regions previously detected. The condition for determining whether an artifact was removed is $\left(\left|A\cap B\right|=0\right)\wedge\left(\left|B\right|\neq 0\right)$. We add a new condition $\left(\left|B\right|\neq 0\right)$ to prevent metrics gaining an increase in this score by introducing new artifacts on previously clear images.
  • $\text{Add}_{pix}$. Pixel-wise addition rate; represents mean percentage of pixels in new artifact regions that resulted from fine-tuning: $\left|A\cap\overline{B}\right|/\left|\overline{B}\right|$.
  • $\text{Rem}_{pix}$. Pixel-wise removal rate; represents average number of pixels that were previously classified as artifacts and that go undetected after fine-tuning: $\left|\overline{A}\cap B\right|/\left|B\right|$.
We report metrics after five epochs of fine-tuning. This is sufficient to saturate the training; Figure 11 shows metric evolution for ten fine-tuning epochs.
We mentioned that we dilated the masks, which improved the fine-tuning quality. We observed that the fine-tuning artificial GT image obtained by masking the upscaled image with the detected artifact mask sometimes produces a poor quality image. In cases when the artifact mask is too tight or noisy, some parts of the artifact region are still present on artificial GT, which may decrease the quality of fine-tuned models. Both of these cases can be solved by dilating the artifact mask before constructing the artificial GT.
Table 7 shows fine-tuning results for all metrics when using dilation (this part of the table matches Table 6 from the main paper), and when using unmodified artifact masks. The proposed dilation step improves the results across all scores and metrics.
Finally, Table 8 provides absolute values for IoU and for the fraction of pixels detected as artifacts, before and after the fine-tuning procedure. $\Delta\text{IoU}$ in the main tables corresponds to the difference between ”IoU before” and ”IoU after”.
Appendix C Subjective quality of images with artifacts
Table 9: Correlation between subjective scores and (1p)(1-p), where $p$ is the prominence score. We calculated subjective scores independently for each group of images that share the same input low-resolution image.
Group Pearson Spearman
image1 1.00 1.00
image2 1.00 0.95
image3 0.93 0.74
image4 0.84 0.95
image5 0.76 0.40
image6 0.47 0.00
image7 0.31 0.20
image8 0.23 0.40
image9 0.22 0.32
image10 0.21 0.40
image11 -0.63 -0.80
image12 -0.76 -0.95
When a super-resolved image contains an artifact, its subjective quality typically decreases, as reflected in assessor scores. This observation shows the importance of artifact detection, since artifact-free images tend to be more visually pleasant than those with artifacts.
To validate this claim, we conducted a side-by-side subjective comparison using images from our dataset. Participants viewed pairs of super-resolved images and were asked to select from each pair the one they preferred. A total of 842 people took part in our subjective study. Calculation of the final scores used the Bradley-Terry model and included 16,720 votes.
We then analyzed the correlation between these subjective scores and (1p)(1-p), where $p$ is the prominence score of the artifact on the image. As Table 9 shows, the correlation was positive in most cases, supporting our hypothesis that artifact-free images tend to receive higher subjective scores.
Appendix D Artifact examples and failure cases
Figure 12: Example artifacts detected by the proposed method. (a): low-resolution input image; (b): target SR result with annotated output artifact mask; (c): artifact prominence heatmap predicted by our method; (d): our input features described in Sec. 4.1, top to bottom: DISTS, bd_jup, ssm_jup.
Figure 13: Example false detections by the proposed method. (a): low-resolution input image; (b): target SR result with annotated output artifact mask; (c): artifact prominence heatmap predicted by our method; (d): our input features described in Sec. 4.1, top to bottom: DISTS, bd_jup, ssm_jup.
Figure 14: Example false detections by the proposed method due to inaccurate restoration from pseudo-GT lightweight SR (RLFN). Rightmost column shows that the false detection disappears when an accurate restoration is used as reference instead of RLFN.
Figure 12 shows examples of prominent artifacts detected by our proposed method across various SR models (Yu et al., 2024; Wang et al., 2021b; 2024a; Liang et al., 2021; Wang et al., 2021a). Each example is annotated with the binary artifact mask and subjective prominence.
Figure 13 shows examples of false detections by our proposed method across SR models (Chen et al., 2023; Hsu et al., 2024b; Yue et al., 2023; Wang et al., 2024b; Cai et al., 2019). We observed the following failure cases:
  • Distortions on natural, unstructured objects, like ground, grass, or trees, that are not very prominent to human observers.
  • Accurate restoration of fine textures such as fur, nylon, or mesh grille. False detections can happen on these when the lightweight SR (in our case, RLFN (Kong et al., 2022)) fails to produce a sharp upscaling of the texture, leading the metrics to see a discrepancy to the target SR and mark it as an artifact. Using an accurately-restored reference removes those false detections as Figure 14 shows.
Existing methods also suffer from these failure cases; indeed, they account for most of the low-prominence detections from our subjective evaluation described in Section 5.3.
Appendix E Subjective evaluation on additional SR datasets
Table 10: Crowd-sourced prominence across SR models on 6 datasets.
SR Type Masks Found Mean Prominence Conf. Masks Found
SeeSR Diffusion 61 7.15% 0
ResShift Diffusion 60 7.20% 0
SinSR Diffusion 60 9.19% 2
HAT-L Transformer 50 9.60% 1
DRCT Transformer 50 13.33% 0
SwinIR Transformer 60 17.12% 4
RealESRGAN CNN 60 17.46% 5
SUPIR Diffusion 62 17.56% 6
Table 11: Crowd-sourced prominence across artifact detection methods on 6 datasets.
Method Masks Found Mean Prominence Conf. Masks Found Comb. Score
ssm_jup (t=0.15) 80 7.42% 1 0.07
bd_jup (t=0.1) 80 7.34% 2 0.15
LDL (t=0.005) 80 9.46% 2 0.19
DeSRA 74 12.21% 3 0.37
Ours (t=0.3) 70 17.85% 8 1.43
DISTS (t=0.25) 76 18.03% 9 1.62
We conduct an additional subjective evaluation on 6 widely known image datasets (Martin et al., 2001; Wang et al., 2018; Dong & Loy, 2016; Bevilacqua et al., 2012; Yang et al., 2010), following the setup described in Section 5.3. In total, this evaluation used 420 source images, each processed by 8 SR models.
Tables 11 and 11 show the results, grouped by SR models and by artifact detection methods, respectively. Interestingly, SR models show much better artifact robustness than in our main comparison in Section 5.3, likely because these datasets are commonly used for SR training and evaluation. Our proposed method falls one confident artifact short of DISTS, but otherwise shows competitive results.
Appendix F F1-score bounds analysis
The theoretical bounds of F1-score are [-0.3, 0.7], governed by the penalty term $\kappa$ = 0.3. The practical range of scores achieved by any non-trivial method depends on the dataset’s GT prominence distribution. Our dataset exhibits the following prominence statistics: mean = 0.40, std. dev. = 0.30, with 53% of masks having a value >>0.3. Compared to the DeSRA dataset (mean = 0.49, std. dev. = 0.26, 71% masks >>0.3), our labels are skewed toward less prominent regions. This difference in dataset bias results in generally higher absolute metric scores on DeSRA.
Appendix G Objective evaluation on the learning-based compression task
Table 12: Results on the full JPEG AI edge artifact dataset.
Method $Prec^{pr}$ $Rec^{pr}$ F1-score PR-AUC
DISTS (t=0.25) 0.0655 0.1783 0.0958 0.0331
[…]
ssm_jup (t=0.2) 0.0784 0.1083 0.0910 0.0324
DeSRA 0.0784 0.0518 0.0623 0.0218
Ours (t=0.15) 0.0742 0.1907 0.1068 0.0295
Ours (t=0.3) 0.0835 0.0883 0.0858
Table 13: Results on the prominent subset of the JPEG AI edge artifact dataset.
Method Precision Recall Prec * Rec IoU PR-AUC
DISTS (t=0.25) 0.0779 0.2339 0.0182 0.0699 0.0444
bd_jup (t=0.1) 0.0467 0.6765 0.0316 0.0912 0.0554
ssm_jup (t=0.2) 0.1051 0.1322 0.0139 0.0584 0.0479
DeSRA 0.1145 0.0481 0.0055 0.0313 0.0414
[…]
Ours (t=0.3) 0.1179 0.0951 0.0112 0.0520
Learning-based image compression has seen a lot of research attention recently, especially as efforts focused on finalizing the JPEG AI compression standard (Ascenso et al., 2023). JPEG AI promises considerable bitrate savings compared to traditional compression. Unfortunately, as Tsereh et al. (2024) show, JPEG AI is also susceptible to the neural artifacts problem, not dissimilar to learning-based super-resolution.
In order to evaluate the transferability of our promised method to other domains, we conduct an objective evaluation on the JPEG AI edge artifact examples dataset collected by Tsereh et al., following our methodology described in Section 5.1. Tables 13 and 13 show the results.
Our proposed method achieves the best F1-score on the full set, and second-best IoU and PR-AUC on the prominent subset, indicating good transferability across domains. We expect the performance to increase further if the proposed method was fine-tuned on artifact examples specific to JPEG AI.
Appendix H Additional experiments
Table 14: Ablation results on the full proposed dataset.
Original HR SPAN RLFN
Method $Prec^{pr}$ $Rec^{pr}$ F1-score PR-AUC $Prec^{pr}$ $Rec^{pr}$ F1-score PR-AUC $Prec^{pr}$ $Rec^{pr}$ F1-score PR-AUC
[…]
w/o bd_jup (t=0.45) 0.0621 0.0485 0.0545 0.0011 0.0454 0.0267 0.0336 0.0008 0.0456 0.0254 0.0327 0.0008
[…]
w/o DISTS (t=0.2) 0.0147 0.0232 0.0180 0.0006 0.0177 0.0209 0.0191 0.0010 0.0095 0.0110 0.0102 0.0009
HR+SPAN+RLFN (t=0.35) 0.0604 0.0470 0.0528 0.0013 0.0588 0.0330 0.0423 0.0022 0.0628 0.0312 0.0417 0.0028
[…]
Just RLFN (t=0.4) 0.0273 0.0322 0.0296 0.0003 0.0885 0.0502 0.0641 0.0013 0.1241 0.0493 0.0706 0.0031
GT-area-only (t=0.35) 0.0224 0.0243 0.0233 0.0107 0.0130 0.0096 0.0110 0.0070 0.0130 0.0090 0.0106 0.0054
Weighted loss (t=0.4) 0.0775 0.0413 0.0539 0.0094 0.0485 0.0195 0.0278 0.0061 0.0539 0.0201 0.0293 0.0061
[…]
Table 15: Ablation results on the full DeSRA dataset.
MSE-SR SPAN RLFN
Method $Prec^{pr}$ $Rec^{pr}$ F1-score PR-AUC $Prec^{pr}$ $Rec^{pr}$ F1-score PR-AUC $Prec^{pr}$ $Rec^{pr}$ F1-score PR-AUC
[…]
w/o bd_jup (t=0.45) 0.1936 0.1846 0.1890 0.0603 0.1059 0.2086 0.1405 0.0375 0.1871 0.1867 0.1869 0.0589
w/o ssm_jup(t=0.25) 0.1718 0.1557 0.1634 0.0467 0.0912 0.1898 0.1232 0.0242 0.1716 0.1606 0.1660 0.0488
w/o DISTS (t=0.2) 0.1308 0.1970 0.1572 0.0346 0.0814 0.2087 0.1171 0.0188 0.1266 0.1986 0.1547 0.0349
[…]
Just RLFN (t=0.4) 0.1415 0.1724 0.1554 0.0405 0.0604 0.1975 0.0925 0.0167 0.1416 0.1755 0.1568 0.0394
GT-area-only (t=0.35) 0.1870 0.1778 0.1823 0.0500 0.1355 0.1922 0.1589 0.0372 0.1781 0.1818 0.1799 0.0521
[…]
Ours (t=0.15) 0.1565 0.2063 0.1780 0.0616 0.0849 0.2262 0.1235 0.0398 0.1500 0.2064 0.1737 0.0605
Ours (t=0.3) 0.2170 0.1701 0.1907 0.1271 0.1955 0.1540 0.2117 0.1727 0.1902
Table 16: Ablation results on the prominent subset of the proposed dataset.
Original HR SPAN RLFN
Method Precision Recall Prec * Rec IoU PR-AUC Precision Recall Prec * Rec IoU PR-AUC Precision Recall Prec * Rec IoU PR-AUC
CNN (t=0.45) 0.4560 0.4836 0.2206 0.3510 0.4495 0.5263 0.320 0.1684 0.2375 0.4034 0.5320 0.3127 0.1664 0.2338 0.4072
[…]
w/o ssm_jup (t=0.25) 0.3454 0.6582 0.2273 0.3686 0.3762 0.4907 0.3527 0.1731 0.2666 0.3807 0.5174 0.2945 0.1524 0.2507 0.3780
w/o DISTS (t=0.2) 0.1693 0.7709 0.1305 0.2920 0.1717 0.1935 0.5818 0.1126 0.2471 0.2055 0.2108 0.5418 0.1142 0.2511 0.2065
[…]
Just RLFN (t=0.4) 0.2260 0.6509 0.1471 0.3028 0.1202 0.3456 0.4982 0.1722 0.2881 0.2574 0.4124 0.3927 0.1619 0.2833 0.3077
GT-area-only (t=0.35) 0.2925 0.7018 0.2053 0.3314 0.3426 0.3524 0.4073 0.1435 0.2567 0.2793 0.3556 0.4073 0.1448 0.2452 0.2903
[…]
Ours (t=0.15) 0.2447 0.8145 0.1993 0.3639 0.4756 0.3709 0.5636 0.2090 0.3018 0.3931 0.3890 0.5273 0.2051 0.2903 0.3829
Ours (t=0.3) 0.4357 0.5745 0.2503 0.3669 0.5209 0.3345 0.1743 0.2357 0.5138 0.3418 0.1756 0.2311
Table 17: Ablation results on the prominent subset of the DeSRA dataset.
MSE-SR SPAN RLFN
Method Precision Recall Prec * Rec IoU PR-AUC Precision Recall Prec * Rec IoU PR-AUC Precision Recall Prec * Rec IoU PR-AUC
[…]
w/o ssm_jup (t=0.25) 0.4964 0.5682 0.2821 0.4433 0.5226 0.2036 0.7352 0.1497 0.3712 0.2748 0.4961 0.6083 0.3018 0.4603 0.5426
w/o DISTS (t=0.2) 0.2165 0.7207 0.1560 0.4428 0.2688 0.1296 0.7753 0.1005 0.3436 0.1644 0.2017 0.7335 0.1480 0.4421 0.2619
[…]
HR+RLFN (t=0.3) 0.4615 0.7159 0.3304 0.5192 0.5892 0.1547 0.8443 0.1306 0.3924 0.3363 0.4367 0.7528 0.3288 0.5263 0.5984
Just RLFN (t=0.4) 0.3399 0.6806 0.2313 0.4522 0.4292 0.1129 0.8138 0.0919 0.2946 0.1636 0.3216 0.7191 0.2312 0.4620 0.4355
GT-area-only (t=0.35) 0.4691 0.6019 0.2824 0.4697 0.3875 0.2741 0.6822 0.1870 0.4243 0.2490 0.4282 0.6453 0.2763 0.4756 0.3939
[…]
Ours (t=0.15) 0.4121 0.7961 0.3281 0.5420 0.6104 0.1633 0.8973 0.1465 0.4010 0.3873 0.3353 0.8427 0.2825 0.5301 0.6030
[…]
This appendix describes our additional experiments and ablation studies. For all experiments, we measure objective metrics as we described in Section 5.1 of the main paper. Tables 14 and 16 show the results on our proposed dataset, and Tables 15 and 17 show the results on the DeSRA (Xie et al., 2023a) dataset.
The following sections describe these model variations in detail.
H.1 Input-feature ablation
As described in Section 4.2, our proposed method takes as input three features: DISTS, bd_jup, and ssm_jup. We train and evaluate three variations of our method, excluding each of these features. These variants are marked “w/o [feature]” in the results tables.
Removing DISTS and ssm_jup results in the heaviest performance drops to our method. This is consistent with the high scores of DISTS alone in our evaluations: this metric is quite capable for detecting artifacts. Removing bd_jup, on the other hand, gives very similar performance to our full proposed method on most evaluations. However, the proposed method still generally outperforms this variant.
H.2 Training on different pseudo-GT
We mention in Section 5.2 that we use RLFN as the pseudo-GT input for all experiments in the paper, unless noted otherwise, since it showed better performance compared to SPAN. It makes sense then that we should also train our proposed method with RLFN passed as the pseudo-GT input. However, this was not the case. We trained our final method using the original high-resolution input, despite evaluating it with the RLFN pseudo-GT.
The results tables show scores for our proposed method’s checkpoints trained using RLFN as pseudo-GT (denoted “Just RLFN”), as well as original high-resolution mixed with RLFN (denoted “HR+RLFN”), and original high-resolution mixed with RLFN and SPAN (denoted “HR+SPAN+RLFN”).
We observe that training on pseudo-GT does tend to improve the results on our proposed dataset when using pseudo-GT for evaluation. However, when evaluating on the DeSRA dataset, training on the original high-resolution frames results in the best method performance, regardless of the pseudo-GT used during evaluation. We hypothesize that this may be because images in the DeSRA dataset have a higher resolution compared to our proposed dataset, making it easy for lightweight SRs to produce high quality pseudo-GT, which appears more similar to the original high-resolution frames.
H.3 GT-area-only training
Our proposed dataset, described in Section 3, includes binary artifact masks and corresponding subjective prominence annotations. During training, our loss function, described in Section 5, moves the model towards predicting the subjective prominence value for pixels inside the binary artifact mask, and 0 (no artifact) outside the binary artifact mask.
However, we considered that this may not be entirely correct. During subjective annotation described in Section 3.1, we ask participants only if they see distortions inside the area denoted by the binary artifact mask. We dim the image to direct the participants’ attention towards the masked area, and away from the other parts of the image. If the binary artifact mask missed an artifact elsewhere on the image, then the participants aren’t expected to see and rank it. Effectively, we may not know the accurate artifact prominence value for regions outside the binary artifact mask, even though in practice those regions do not contain artifacts.
We conducted an experiment to account for this in training by disabling the loss component responsible for the region outside the binary artifact mask. This way, our loss function only considered the pixels within the artifact mask—for which we know the ground-truth subjective prominence value. Given our dataset’s bias towards low-prominence samples, the model should still be able to learn when to predict the absence of an artifact.
The model from this training run is labeled “GT-area-only” in the results tables. We found that it failed to match the performance of the proposed model with the full loss function. Our hypothesis for this outcome is that the regions outside the binary artifact mask indeed contain no artifacts in most cases, and their loss function component helps the model learn to better localize artifacts on an image.
H.4 Weighting loss to normalize training-set classes
As we pointed out in Section 3.3, our proposed dataset has a bias towards low-prominence samples. In this experiment, we tried to account for this bias in training by weighting the loss function according to the number of samples with the given prominence (inversely proportional to the histogram on Figure 4).
The model from this training run is labeled “Weighted loss” in the results tables. It shows similar performance to the proposed model with uniformly weighted loss, and even overtakes it in specific scenarios, but the proposed model has better results overall. During training, we observed very similar validation curves between weighted and non-weighted loss.
H.5 Variance across training runs
Table 18: Statistics across 4 proposed model checkpoints for full DeSRA and proposed datasets
F1-score PR-AUC
Dataset Reference Input Threshold Mean Std Mean Std
Ours Original HR 0.15 0.0273 0.0086 0.0118 0.0009
0.30 0.0530 0.0126
RLFN 0.15 0.0285 0.0023 0.0081 0.0004
0.30 0.0346 0.0074
DeSRA MSE-SR 0.15 0.1644 0.0154 0.0608 0.0010
0.30 0.1866 0.0050
RLFN 0.15 0.1586 0.0161 0.0595 0.0012
0.30 0.1837 0.0072
We trained our model across four random seeds. Table 18 demonstrates strong stability across seeds, as indicated by consistently low standard deviations relative to the mean values for both performance metrics. The PR AUC exhibits good stability, with fairly low standard deviations (0.0004–0.0012) across all dataset and ground-truth configurations. Similarly, the F1-score shows good reproducibility, on the both datasets. Overall, the low variance confirms that the reported mean performance metrics are highly reproducible and not artifacts of a single model initialization.
H.6 Using a CNN instead of a per-pixel MLP
As we mentioned in Section 4.2, our proposed multilayer perceptron processes features for every pixel of the input image individually, which does not preclude it from using wider context, because our input features (DISTS, bd_jup, ssm_jup) themselves use wider context during computation. However, it is a reasonable assumption that the artifact prominence regressor itself may be able to extract additional useful contextual information from surrounding pixels, as the input features were not trained specifically for the artifact detection task.
We conducted an experiment replacing the multilayer perceptron with a small CNN. It consisted of five consecutive 3×3 convolution + ReLU + layer normalization residual blocks, each with 8 depth channels.
This model is labeled “CNN” in the results tables. We tried both passing just the original three features as input, and the original three features together with the normalized R, G, and B color channels of the input image. Both of these variants netted similar results, which failed to improve upon our proposed multilayer perceptron model.
H.7 Using a random forest instead of MLP
Table 19: $F1{\text{-}}score^{pr}$ metric for random-forest checkpoints, DeSRA, and proposed method.
Number of estimators Max depth
2 4 6 8 10 12
8 0.0250 0.0259 0.0304 0.0345 0.0362 0.0361
16 0.0250 0.0259 0.0304 0.0356 0.0366 0.0364
32 0.0250 0.0261 0.0299 0.0353 0.0365 0.0364
[…]
Proposed MLP method (t=0.15) 0.0355
Proposed MLP method (t=0.3) 0.0559
We trained several random-forest models on the features described in Section 4.1 using different sets of hyperparameters: the number of estimators varied between 8, 16, and 32, and the maximum tree depth was 2, 4, 6, 8, 10, or 12. We used the random-forest implementation from the XGBoost library. All other hyperparameters except n_estimators and max_depth remain their default values. Empirically, we chose 0.05 as the best threshold for all trained random-forest models.
Our comparison used the test set and the $F1{\text{-}}score^{pr}$ metric. Table 19 shows the prediction scores of the trained models compared with our proposed method. Note that the random-forest architecture fails to achieve higher quality than DeSRA or our multilayer-perceptron architecture.
The experiments also investigated other training-data-preparation approaches, but they failed to exhibit strong increases in the quality of artifact-prominence prediction. In particular, neither block averaging of ground-truth labels nor exclusion from the training of the regions that had not been labeled as containing artifacts during the subjective comparison.
H.8 Filtering heatmaps through semantic segmentation
Similarly to Xie et al. (2023a), we noticed that SR-artifact prominence is related to the affected object. Viewers hardly notice artifacts in objects such as grass, leaves, dirt, and soil. To confirm this hypothesis, we used the SAN (Xu et al., 2023) semantic segmentation method to cancel artifact predictions on pixels of certain semantic classes.
To select classes for exclusion, we analyzed the training set and calculated the average prominence for classes that occur in the artifact regions, then selected those classes with an average prominence less than 0.3. From these classes we manually removed those that define objects with potentially prominent artifacts—for example, “mouse,” “umbrella,” “cake,” and so on. The remaining classes were used to exclude such objects from artifact prediction.
This procedure improves the $F1{\text{-}}score^{pr}$ metric (Table 20), which considers the artifact’s prominence. Accordingly, excluding special classes with artifacts that are difficult to notice avoids penalizing the methods. However, this procedure fails to improve the results on the DeSRA dataset annotated in lab (Table 21), perhaps because the artifacts on organic matter were missed during annotation due to low prominence. Next, in a scenario with highly prominent test-set sampling, excluding classes only degrades IoU (Table 22, Table 23), because in this case all artifacts in the dataset are already prominent.
Considering these results, we have decided to omit this filtering step from our final proposed method, saving runtime and resources.
Table 20: SAN comparison on the full proposed dataset.
Original HR SPAN RLFN
Method $Prec^{pr}$ $Rec^{pr}$ F1-score PR-AUC $Prec^{pr}$ $Rec^{pr}$ F1-score PR-AUC $Prec^{pr}$ $Rec^{pr}$ F1-score PR-AUC
LDL (t=0.005) 0.0995 0.0159 0.0275 0.0039 0.0325 0.0141 0.0197 0.0034 0.0054 0.0026 0.0035 0.0007
SSIM (t=0.55) 0.0120 0.0169 0.0140 0.0022 0.0538 0.0269 0.0359 0.0083 0.0541 0.0252 0.0344 0.0054
LPIPS (t=0.25) 0.0680 0.0238 0.0352 0.0049 0.1094 0.0258 0.0418 0.0042 0.0893 0.0228 0.0364 0.0043
ERQA (t=0.55) 0.0024 0.0032 0.0028 0.0001 -0.0163 -0.0119 -0.0137 -0.0002 -0.0262 -0.0155 -0.0195 -0.0004
PAL4Inpaint (bin., no-ref) 0.0208 0.0081 0.0117 N/A 0.0208 0.0081 0.0117 N/A 0.0208 0.0081 0.0117 N/A
PAL4VST (bin., no-ref) 0.0464 0.0033 0.0062 N/A 0.0464 0.0033 0.0062 N/A 0.0464 0.0033 0.0062 N/A
[…]
bd_jup (t=0.1) 0.0032 0.0064 0.0043 0.0028 0.0088 0.0131 0.0105 0.0013 0.0063 0.0091 0.0074 0.0017
ssm_jup (t=0.2) 0.0266 0.0237 0.0251 0.0012 0.0165 0.0127 0.0144 0.0011 0.0209 0.0158 0.0180 0.0009
[…]
Ours (t=0.3) 0.0762 0.0441 0.0559 0.0503 0.0224 0.0310 0.0581 0.0235 0.0334
LDL + SAN (t=0.005) 0.1458 0.0171 0.0307 0.0055 0.1221 0.0305 0.0488 0.0104 0.1128 0.0271 0.0438 0.0081
SSIM + SAN (t=0.55) 0.0623 0.0532 0.0574 0.0098 0.1454 0.0421 0.0652 0.0142 0.1406 0.0384 0.0604 0.0116
LPIPS + SAN (t=0.25) 0.0766 0.0221 0.0344 0.0061 0.1560 0.0244 0.0421 0.0074 0.1416 0.0232 0.0399 0.0077
ERQA + SAN (t=0.55) 0.0579 0.0471 0.0520 0.0021 0.0559 0.0231 0.0327 0.0017 0.0568 0.0181 0.0275 0.0016
PAL4Inpaint + SAN (bin., no-ref) 0.0268 0.0091 0.0136 N/A 0.0268 0.0091 0.0136 N/A 0.0268 0.0091 0.0136 N/A
PAL4VST + SAN (bin., no-ref) 0.0419 0.0028 0.0053 N/A 0.0419 0.0028 0.0053 N/A 0.0419 0.0028 0.0053 N/A
[…]
bd_jup + SAN (t=0.1) 0.0393 0.0522 0.0448 0.0069 0.0571 0.0534 0.0552 0.0063 0.0552 0.0500 0.0525 0.0061
ssm_jup + SAN (t=0.2) 0.0810 0.0460 0.0587 0.0057 0.0824 0.0386 0.0526 0.0062 0.0867 0.0397 0.0544 0.0056
[…]
Ours + SAN (t=0.3) 0.1274 0.0542 0.0761 0.1396 0.0380 0.0598 0.1451 0.0369 0.0588
Table 21: SAN comparison on the full DeSRA dataset.
MSE-SR SPAN RLFN
Method $Prec^{pr}$ $Rec^{pr}$ F1-score PR-AUC $Prec^{pr}$ $Rec^{pr}$ F1-score PR-AUC $Prec^{pr}$ $Rec^{pr}$ F1-score PR-AUC
LDL (t=0.005) 0.2320 0.1305 0.1670 0.0518 0.1615 0.1621 0.1618 0.0486 0.2339 0.1242 0.1622 0.0503
[…]
LPIPS (t=0.25) 0.1543 0.1268 0.1392 0.0349 0.1304 0.1445 0.1371 0.0300 0.1591 0.1351 0.1462 0.0389
ERQA (t=0.55) 0.0711 0.0275 0.0396 0.0026 0.0396 0.0590 0.0474 0.0017 0.0728 0.0275 0.0399 0.0025
PAL4Inpaint (bin., no-ref) 0.0543 0.0693 0.0609 N/A 0.0543 0.0693 0.0609 N/A 0.0543 0.0693 0.0609 N/A
PAL4VST (bin., no-ref) 0.0243 0.0030 0.0054 N/A 0.0243 0.0030 0.0054 N/A 0.0243 0.0030 0.0054 N/A
DISTS (t=0.25) 0.1478 0.1813 0.1628 0.0376 0.0717 0.2115 0.1071 0.0213 0.1470 0.1847 0.1637 0.0457
bd_jup (t=0.1) 0.0825 0.2043 0.1175 0.0230 0.0585 0.2153 0.0920 0.0135 0.0825 0.2079 0.1181 0.0244
ssm_jup (t=0.2) 0.1900 0.1655 0.1769 0.0377 0.1170 0.1825 0.1426 0.0250 0.1717 0.1663 0.1690 0.0346
[…]
Ours (t=0.15) 0.1565 0.2063 0.1780 0.0616 0.0849 0.2262 0.1235 0.0398 0.1500 0.2064 0.1737 0.0605
[…]
SSIM + SAN (t=0.55) 0.2238 0.1447 0.1758 0.0513 0.1386 0.1612 0.1491 0.0327 0.2119 0.1461 0.1729 0.0500
LPIPS + SAN (t=0.25) 0.1806 0.1035 0.1316 0.0328 0.1491 0.1211 0.1336 0.0259 0.1854 0.1100 0.1380 0.0342
ERQA + SAN (t=0.55) 0.0905 0.0254 0.0396 0.0030 0.0463 0.0506 0.0484 0.0018 0.0919 0.0257 0.0402 0.0029
PAL4Inpaint + SAN (bin., no-ref) 0.0556 0.0547 0.0552 N/A 0.0556 0.0547 0.0552 N/A 0.0556 0.0547 0.0552 N/A
PAL4VST + SAN (bin., no-ref) 0.0283 0.0028 0.0050 N/A 0.0283 0.0028 0.0050 N/A 0.0283 0.0028 0.0050 N/A
DISTS + SAN (t=0.25) 0.1748 0.1535 0.1634 0.0376 0.0820 0.1774 0.1121 0.0171 0.1742 0.1562 0.1647 0.0374
bd_jup + SAN (t=0.1) 0.0973 0.1737 0.1247 0.0210 0.0676 0.1825 0.0987 0.0124 0.0970 0.1763 0.1251 0.0219
ssm_jup + SAN (t=0.2) 0.2065 0.1393 0.1663 0.0321 0.1290 0.1522 0.1396 0.0216 0.1887 0.1399 0.1607 0.0297
DeSRA + SAN (t=0.3) 0.2460 0.1262 0.1668 0.0542 0.1435 0.1084 0.1235 0.0261 0.2340 0.1231 0.1614 0.0513
Ours + SAN (t=0.15) 0.1782 0.1739 0.1760 0.0563 0.0951 0.1887 0.1265 0.0367 0.1723 0.1743 0.1733 0.0549
Ours + SAN (t=0.3) 0.2384 0.1416 0.1777 0.1390 0.1619 0.1496 0.2324 0.1437 0.1776
Table 22: SAN comparison on the prominent subset of the proposed dataset.
Original HR SPAN RLFN
Method Precision Recall Prec * Rec IoU PR-AUC Precision Recall Prec * Rec IoU PR-AUC Precision Recall Prec * Rec IoU PR-AUC
LDL (t=0.005) 0.3684 0.1418 0.0522 0.1043 0.1387 0.3764 0.3055 0.1150 0.1788 0.3110 0.3942 0.3091 0.1218 0.1779 0.3361
[…]
LPIPS (t=0.25) 0.5876 0.2473 0.1453 0.2621 0.3860 0.4983 0.1418 0.0707 0.1340 0.3044 0.4393 0.1600 0.0703 0.1324 0.2971
ERQA (t=0.55) 0.1957 0.6218 0.1217 0.2495 0.1220 0.2300 0.3236 0.0744 0.1670 0.1052 0.2494 0.2327 0.0581 0.1492 0.1052
PAL4Inpaint (bin., no-ref) 0.1833 0.1164 0.0213 0.0753 N/A 0.1833 0.1164 0.0213 0.0753 N/A 0.1833 0.1164 0.0213 0.0753 N/A
PAL4VST (bin., no-ref) 0.2682 0.0327 0.0088 0.0463 N/A 0.2682 0.0327 0.0088 0.0463 N/A 0.2682 0.0327 0.0088 0.0463 N/A
[…]
bd_jup (t=0.1) 0.1810 0.8364 0.1514 0.2843 0.3311 0.1800 0.6945 0.1250 0.2475 0.2342 0.1983 0.6509 0.1291 0.2434 0.2275
ssm_jup (t=0.2) 0.2350 0.4909 0.1153 0.2368 0.2127 0.2383 0.4255 0.1014 0.2133 0.2273 0.2689 0.4582 0.1232 0.2221 0.2411
DeSRA 0.4791 0.2764 0.1324 0.2560 0.3173 0.6976 0.0982 0.0685 0.1296 0.3358 0.7040 0.0836 0.0589 0.1205 0.3025
[…]
Ours (t=0.3) 0.4357 0.5745 0.2503 0.3669 0.5209 0.3345 0.1743 0.2357 0.5138 0.3418 0.1756 0.2311
LDL + SAN (t=0.005) 0.3927 0.1236 0.0486 0.0964 0.1296 0.4220 0.2764 0.1166 0.1692 0.3086 0.4260 0.2800 0.1193 0.1680 0.3179
SSIM + SAN (t=0.55) 0.3164 0.6436 0.2036 0.3476 0.3759 0.5175 0.3164 0.1637 0.2500 0.3608 0.5315 0.2800 0.1488 0.2294 0.3552
LPIPS + SAN (t=0.25) 0.6108 0.2145 0.1310 0.2417 0.3694 0.5437 0.1273 0.0692 0.1224 0.2750 0.4918 0.1382 0.0680 0.1189 0.2748
ERQA + SAN (t=0.55) 0.1979 0.5927 0.1173 0.2459 0.1194 0.2497 0.3091 0.0772 0.1605 0.1067 0.2745 0.2182 0.0599 0.1428 0.1055
PAL4Inpaint + SAN (bin., no-ref) 0.1780 0.1127 0.0201 0.0706 N/A 0.1780 0.1127 0.0201 0.0706 N/A 0.1780 0.1127 0.0201 0.0706 N/A
PAL4VST + SAN (bin., no-ref) 0.2591 0.0291 0.0075 0.0402 N/A 0.2591 0.0291 0.0075 0.0402 N/A 0.2591 0.0291 0.0075 0.0402 N/A
DISTS + SAN (t=0.25) 0.3167 0.5600 0.1773 0.3339 0.3491 0.4842 0.3673 0.1778 0.2570 0.3217 0.5069 0.3345 0.1696 0.2525 0.3260
bd_jup + SAN (t=0.1) 0.1722 0.7600 0.1309 0.2744 0.3243 0.1821 0.6218 0.1132 0.2360 0.2226 0.1990 0.5818 0.1158 0.2329 0.2112
ssm_jup + SAN (t=0.2) 0.2490 0.4473 0.1114 0.2247 0.2026 0.2454 0.3855 0.0946 0.2007 0.2072 0.2830 0.4145 0.1173 0.2101 0.2319
DeSRA + SAN (t=0.3) 0.5160 0.2545 0.1313 0.2419 0.3041 0.7193 0.0909 0.0654 0.1225 0.3120 0.7069 0.0764 0.0540 0.1128 0.2763
[…]
Table 23: SAN comparison on the prominent subset of the DeSRA dataset.
MSE-SR SPAN RLFN
Method Precision Recall Prec * Rec IoU PR-AUC Precision Recall Prec * Rec IoU PR-AUC Precision Recall Prec * Rec IoU PR-AUC
LDL (t=0.005) 0.5929 0.3772 0.2237 0.3724 0.4687 0.2977 0.5409 0.1610 0.3896 0.3418 0.5744 0.3467 0.1992 0.3443 0.4183
[…]
LPIPS (t=0.25) 0.5152 0.4398 0.2266 0.3759 0.4488 0.3385 0.4526 0.1532 0.3450 0.3193 0.5261 0.5072 0.2669 0.4094 0.5014
ERQA (t=0.55) 0.1365 0.0401 0.0055 0.0523 0.0100 0.0618 0.2295 0.0142 0.0684 0.0154 0.1422 0.0337 0.0048 0.0514 0.0087
PAL4Inpaint (bin., no-ref) 0.1184 0.1958 0.0232 0.1139 N/A 0.1184 0.1958 0.0232 0.1139 N/A 0.1184 0.1958 0.0232 0.1139 N/A
PAL4VST (bin., no-ref) 0.0407 0.0177 0.0007 0.0140 N/A 0.0407 0.0177 0.0007 0.0140 N/A 0.0407 0.0177 0.0007 0.0140 N/A
DISTS (t=0.25) 0.3898 0.7400 0.2884 0.4919 0.4408 0.1097 0.8604 0.0944 0.3479 0.2290 0.3801 0.7576 0.2880 0.5016 0.5428
bd_jup (t=0.1) 0.1245 0.8347 0.1039 0.3580 0.1609 0.0919 0.8652 0.0795 0.2798 0.1221 0.1160 0.8555 0.0992 0.3625 0.1773
ssm_jup (t=0.2) 0.4629 0.5120 0.2370 0.4032 0.3889 0.3018 0.6164 0.1860 0.3770 0.2736 0.4087 0.5185 0.2119 0.3930 0.3646
DeSRA 0.6794 0.6228 0.4231 0.5277 0.6928 0.3324 0.4462 0.1483 0.3707 0.2910 0.6366 0.6100 0.3883 0.5082 0.6614
[…]
LDL + SAN (t=0.005) 0.5885 0.3114 0.1833 0.3007 0.3662 0.3039 0.4141 0.1259 0.3047 0.2717 0.5809 0.2921 0.1697 0.2814 0.3322
SSIM + SAN (t=0.55) 0.5145 0.5088 0.2618 0.4092 0.4605 0.2510 0.5955 0.1495 0.3510 0.2765 0.4908 0.5169 0.2537 0.4027 0.4519
LPIPS + SAN (t=0.25) 0.4990 0.3162 0.1578 0.2754 0.3175 0.3184 0.3483 0.1109 0.2588 0.2153 0.5128 0.3708 0.1901 0.3014 0.3433
ERQA + SAN (t=0.55) 0.1534 0.0401 0.0062 0.0481 0.0117 0.0672 0.1990 0.0134 0.0627 0.0134 0.1664 0.0337 0.0056 0.0475 0.0100
PAL4Inpaint + SAN (bin., no-ref) 0.1150 0.1621 0.0186 0.0859 N/A 0.1150 0.1621 0.0186 0.0859 N/A 0.1150 0.1621 0.0186 0.0859 N/A
PAL4VST + SAN (bin., no-ref) 0.0448 0.0177 0.0008 0.0130 N/A 0.0448 0.0177 0.0008 0.0130 N/A 0.0448 0.0177 0.0008 0.0130 N/A
DISTS + SAN (t=0.25) 0.3723 0.5698 0.2121 0.3733 0.3134 0.1052 0.6822 0.0718 0.2653 0.1363 0.3658 0.5811 0.2125 0.3824 0.3219
bd_jup + SAN (t=0.1) 0.1217 0.6613 0.0805 0.2887 0.1253 0.0823 0.6918 0.0569 0.2256 0.0872 0.1118 0.6822 0.0763 0.2936 0.1243
ssm_jup + SAN (t=0.2) 0.4549 0.4125 0.1876 0.3294 0.3165 0.2835 0.4848 0.1374 0.3020 0.2158 0.4032 0.4173 0.1683 0.3252 0.2963
DeSRA + SAN (t=0.3) 0.6685 0.4880 0.3262 0.4059 0.5389 0.3211 0.3451 0.1108 0.2779 0.2211 0.6269 0.4751 0.2978 0.3902 0.5166
Ours + SAN (t=0.15) 0.3955 0.6292 0.2489 0.4222 0.4864 0.1496 0.7143 0.1068 0.3080 0.2981 0.3646 0.6372 0.2323 0.4197 0.4832
Ours + SAN (t=0.3) 0.5810 0.4398 0.2555 0.3758 0.2773 0.5602 0.1554 0.3360 0.5723 0.4655 0.2664 0.3851